OpenStudy (anonymous):

Anyone? When I went to the store, I purchased four items. The following shows the cost of three of the items: $1.50 $3.00 $4.00 The line to the checkout was pretty long, so to quench my boredom I started playing with my pocket calculator while waiting. I found out, to my surprise, that the four prices of the four items I purchased added to the same number as I got when I multiplied the four prices together. What was the price of the fourth item? :))) Medal for the first to answer...

OpenStudy (ja1):

Is this a riddle or a question you need help with?

OpenStudy (anonymous):

A riddle... :)))

OpenStudy (anonymous):

google it XD The three prices add to 1.50 + 3.00 + 4.00 = $8.50. The three prices multiply to 1.50 x 3.00 x 4.00 = $18.00. You might be able to see that adding another $0.50 will take the total to $9.00, and multiplying by another $0.50 will take the product to $9.00 also, but the answer can also be found using algebra. Let's call the unknown price P: 1.50 + 3.00 + 4.00 + P = 1.50 x 3.00 x 4.00 x P 8.50 + P = 18.00 x P 8.50 = 18.00 x P - P 8.50 = 17.00 x P 8.50/17.00 = P P = 0.50.

OpenStudy (aayushi.somani):

i got d same..!!! but by own.. and not google..!!! XD

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

@sarah786 You didn't take the challenge to answer it by yourself, afterall....

OpenStudy (aayushi.somani):

coz u already gave the answer..!! 5 sec before i typed mne soo i had to erase mine..!! @sarah786

OpenStudy (anonymous):

And I was abt to say not to google it...

OpenStudy (anonymous):

lol I don't saying anything . I dont know it is challenge .and you didn't say Google is not allowed XD

OpenStudy (aayushi.somani):

u were quite late @liliegirl .. @sarah786 is very fast..!!!

OpenStudy (anonymous):

haha, true..

OpenStudy (anonymous):

Im not fast . Google is very Fast :D

OpenStudy (aayushi.somani):

nd you prefer using google then your mind..!! coz ur mind is slow.??? @sarah786

OpenStudy (anonymous):

anyways, want some more? and @sarah786 don't you dare google it ... xD @aayushi.somani YEAH RIGHT. CHEERS..

OpenStudy (anonymous):

lol it is . right now . XD ਉ_ਉ

OpenStudy (anonymous):

lol @liliegirl Carry on . I'll not interfere XD

OpenStudy (aayushi.somani):

yeah post some more @liliegirl

OpenStudy (anonymous):

... Man Wrinkle spent one-fourth of his life as a boy, one-eighth as a youth, and one-half as an active man. If Man Wrinkle spent 11 years as an old man, then how many years did he spend as an active man?

OpenStudy (aayushi.somani):

he spent 11 yrs as youth.. 5.5 yrs as a boy and 2.25yrs as an active man..

OpenStudy (aayushi.somani):

m i right @liliegirl and @sarah786 ???

OpenStudy (anonymous):

heyyy!!!!!!!! who mentioned me ..lol I told you Im Out XD

OpenStudy (anonymous):

Cheers, @aayushi.somani

OpenStudy (aayushi.somani):

y i mean m sorry if u felt bad by my comment i was just joking..!!! :P

OpenStudy (aayushi.somani):

thanks @liliegirl

OpenStudy (anonymous):

lol its ok @aayushi.somani . :D

OpenStudy (aayushi.somani):

fine..! @sarah786 .. btw do u have more @liliegirl

OpenStudy (anonymous):

... On a 12-hour digital clock, what is the smallest interval between two times that are palindromic (can be read forwards and backwards as the same number)?

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (aayushi.somani):

21:21 .. i guess.. or 12:12.. may be @liliegirl

OpenStudy (anonymous):

Nope.. wrong... Example of palindromic: 12:21, 9:39 And the question is smallest interval..

OpenStudy (anonymous):

In five minutes, I'll be out, so tell me if you can't figure out the answer... :))))

OpenStudy (anonymous):

http://www.braingle.com/Math.html Bye... Btw, the answer is 2 minutes.. from 9:59 to 10:01 Thanks. Till the next time. :))))